40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Mar 27 20:41:21 2017, rana, gautam, Update, PSL, PMC DAQ assay for feed-forward integration TEK00000.PNG20170327_194931.jpg20170327_204554.jpg
    Reply  Wed Mar 29 11:41:19 2017, gautam, Update, PSL, PMC DAQ assay for feed-forward integration C1ALS_PSLnamespace.pngPMCcheckout.pdfD980352-A-40m_151119.pdf
       Reply  Thu Mar 30 00:16:09 2017, gautam, Update, PSL, PSL NPRO PZT calibration 
          Reply  Mon Apr 3 17:25:13 2017, gautam, Update, PSL, PSL NPRO PZT calibration PDHerr.pdfNPROcalib.pdf
             Reply  Tue Apr 4 17:27:58 2017, rana, Update, PSL, PSL NPRO PZT calibration 
       Reply  Thu Mar 30 18:11:01 2017, rana, Update, PSL, PMC DAQ assay for feed-forward integration 
          Reply  Mon Apr 3 23:07:09 2017, gautam, Update, PSL, PMC DAQ assay for feed-forward integration 
             Reply  Tue Apr 11 00:38:37 2017, gautam, Update, PSL, PMC demod moved off servo board PMC_Ctrl_spec.pdf
                Reply  Wed Apr 12 00:36:53 2017, gautam, Update, PSL, PMC demod moved off servo board PMC_digitalSpec.pdfPMC_DAQ_spectra.pdf
                   Reply  Tue Apr 18 01:01:03 2017, gautam, Update, PSL, PMC OLTF measured, DAQ channels calibrated PMCspectra_calibrated.pdfPMC_MEDMs.png
                      Reply  Tue Apr 18 16:10:00 2017, gautam, Update, PSL, PMC OLTF measured, DAQ channels calibrated PMC_OLTF_170418.pdf
                         Reply  Tue Apr 18 23:37:15 2017, rana, Update, PSL, PMC OLTF measured, DAQ channels calibrated 
                         Reply  Wed Apr 19 15:13:30 2017, gautam, Update, PSL, PMC/MCL multicoherence PMC_MCL_multicoherence.pdf
Message ID: 12945     Entry time: Tue Apr 18 16:10:00 2017     In reply to: 12944     Reply to this: 12946   12947
Author: gautam 
Type: Update 
Category: PSL 
Subject: PMC OLTF measured, DAQ channels calibrated 

Here are the details:

  1. PMC OLTF:
    • the procedure used was identical to what Koji describes in this entry.
    • I used the SR785 to take the measurement.
    • MEDM gain slider was at +20dB 
    • I used the two single pin LEMO front panel monitor points to make the measurement. 
    • Mix_out_mon was CH2A, HV_out_mon was CH1A on the SR785
    • A = CH2A/CH1A with the SR785 excitation applied to the EXT_DC single pin LEMO input on the front panel. I used an excitation amplitude of 15mV
    • B = CH2A/CH1A without any excitation
    • Couple of lines of loop algebra tells us that the OLTF is given by the ratio A/B. The plot below lines up fairly well with what Koji measured here, UGF is ~3.3kHz with a phase margin of ~60degrees, and comparable gain margin at ~28kHz. As noted by Koji, the feature at ~8kHz prevents further increase of the servo gain. I've updated the nominal gain on the PMC MEDM screen accordingly... I couldn't figure out how to easily extract Koji's modelled OLTF so I didn't overlay that here... Overlaid is the model OLTF. No great care was taken in analyzing the goodness of the agreement with the model and measurement by looking at residuals etc, except that the feature that was previously at 28.8kHz now seems to have migrated to about 33.5 kHz. I'm not sure what to make of that. 
  2. PMC DAQ calibration:
    • The calibration was done using the swept cavity, the procedure is basically the same as described by Koji in this elog.
    • The procedure was slightly complicated by the fact that I added gain to the AD620 buffers that provide the DAQ signals. So simply sweeping the cavity saturates the AD620 very quickly.
    • To workaround this, I first hooked up the un-amplified single pin LEMO front panel monitor points to the DAQ channels using some of the available BNC-LEMO patch cables.
    • I then did the swept cavity measurement, and recorded the error and control signals fron the single pin LEMO front panel monitor points. Sweep signal was applied to EXT_DC input on front panel.
    • In the nominal DAQ setup however, we have the amplification on the AD620. I measured this amplification factor by hooking up the single pin LEMO monitor point, along with its corresponding AD620 amplified counterpart, to an SR785 and measuring the transfer function. For the PMC_ERR channel, the AD620 gain is ~53.7dB (i.e. approx 484x). For the PMC_CTRL channel, the AD620 gain is ~33.6dB (i.e. approx 48x). These numbers match up well with what I would expect given the resistors I installed on the PMC board between pins 1 and 8 of the AD620. These gains are digitally undone in the corresponding filter modules, FM1.
    • To calibrate the time axis into frequency, I located the zero crossings of the sidebands and equated the interval to 2 x fmod. For the PMC servo, fmod = 35.5MHz. I used ~1Hz triangle wave, 2Vpp to do the sweep. The resulting slope was 1.7026 GHz/s.
    • The linear part of the PDH error signal for the carrier resonance was fitted with a line. It had a slope of 1.5*10^6 cts/s.
    • The round trip length of the PMC cavity was assumed to be 0.4095m as per Koji's previous entry. This allows us to calibrate the swept cavity motion from Hz to m. The number is 1.4534 * 10^-15 m/Hz. I guess we could confirm this by sweeping the cavity with the DC bias slider through the full range of 0-250V, but we only have a slow readback of the PMC reflection (and no readback of the PMC transmission).
    • Putting the last three numbers together, I get the PMC_ERR signal calibration as 1.6496 pm/ct. This is the number in the "cts2m" filter module (FM10).
    • An analogous procedure was done to calibrate the control signal slope: from the sweep, I got 4617 cts/s, which corresponds to 2.7117*10^-6 cts/Hz. Using the FSR to convert into cts/m, I get for PMC_CTRL, 535.96 pm/ct. This is the number in the "cts2m" filter module (FM10).
    • For convenience, I also added "cts2Hz" calibration filters in FM9 in the corresponding filter modules. 

The updated schematic with changes made, along with some pictures, have been uploaded to the DCC page...

Quote:

Quick entry, details to follow in the AM tomorrow.​

 

Attachment 1: PMC_OLTF_170418.pdf  61 kB  Uploaded Wed Apr 19 00:28:32 2017  | Hide | Hide all
PMC_OLTF_170418.pdf
ELOG V3.1.3-